OpenStudy (anonymous):

A = P(1 + i)n A = Future Value P = Present Value i = interest rate per period n = number of compounding period Important Reminders: • Use a banker's year: 1 year = 360 days • Don’t round off an answer until the final step. • Round all final answers to the nearest cent (hundredth) or the nearest tenth of a percent (thousandth). ________________________________________ A. Find the future value of each of the following deposits. 1) P = $345, 10 years, 6% compounded annually

OpenStudy (phi):

in \[ A= P(1+i)^n \] for 1) P = $345, 10 years, 6% compounded annually what is P? what is i? what is n?

OpenStudy (anonymous):

P= Principal cash I= Interest rate N= # of times compounded

OpenStudy (phi):

yes, but from the problem, can you assign them numbers?

OpenStudy (anonymous):

A=$345(1+.06)^10 is that right?

OpenStudy (phi):

yes you can simplify 1+0.06 to 1.06 type 345*(1.06)^10= in the google search window (or use your calculator)

OpenStudy (phi):

i get 617.842455 round to the nearest penny: the digit after .84 is 2 . 2 is less than 5 so round to 0.84 617.84 is the answer.

OpenStudy (anonymous):

yah i got 617.8424553 to.

OpenStudy (anonymous):

THX!!!

OpenStudy (phi):

be sure to round it to the nearest cent.